# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ 4.2 Single-coalition-first action models
A single-coalition-first action model is a tuple $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{AC}$ is a nonempty set of actions.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathsf{out}_{a}:\mathtt{ST}\times\mathtt{JA}_{a}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $a$ such that for all $s\in\mathtt{ST}$, $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathtt{suc}(s)$.
Definition 15 (Outcome functions and availability functions of single-coalition-first action models).
Let $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first action model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\begin{cases}\mathtt{suc}(s)&\text{if $\mathrm{C}=\emptyset$}\\ \bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}&\text{if $\mathrm{C}\neq\emptyset$}\end{cases}$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{av}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$
@@
For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ equals to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
The following result offers two alternative definitions of sing-coalition-first action models.
@@
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
@@
      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
    It is easy to check this set of conditions is equivalent to the set of constraints on single-coalition-first action models given in Definition 14.
@@
      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}\neq\emptyset$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. It suffices to show $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\emptyset}(s,\emptyset)$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $a\in\mathrm{D}$. By (1b), $t\in\mathsf{out}_{a}(s,\sigma_{\mathrm{D}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ $=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
First, we show (3a). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
@@
By (1a) and (1b), it is easy to verify $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Assume ${\overline{\mathrm{C}}}=\emptyset$. Clearly, what we want holds. Assume ${\overline{\mathrm{C}}}\neq\emptyset$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there is $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})$, …, there is $\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$ such that $w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
@@
Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which equals to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
@@
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model indicated in Figure 1 is a counterexample.
@@
Goranko and Jamroga [4] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [1]. It can be verified convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
@@ 4.3 Single-coalition-first neighborhood models
Let $\mathtt{ST}$ be a nonempty set of states and $\Delta_{1},\Delta_{2}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}$. Define $\Delta_{1}\odot\Delta_{2}=\{\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\mid\mathrm{Y}_{1}\in\Delta_{1},\mathrm{Y}_{2}\in\Delta_{2},\text{and }\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset\}$. Let $I$ be a nonempty set of indices. Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected. Note $\bigodot\{\Delta\}=\Delta$.
@@
A single-coalition-first neighborhood model is a tuple $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • $\mathtt{suc}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{ST})$ is a successor function.
  • for all $a\in\mathtt{AG}$, $\mathtt{nei}_{a}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $a$ such that for all $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$.
Definition 17 (Neighborhood functions of single-coalition-first neighborhood models).
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the neighborhood function $\mathtt{nei}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$,
$\mathtt{nei}_{\mathrm{C}}(s)=\begin{cases}\emptyset&\text{if }\mathtt{suc}(s)=\emptyset\\ \{\mathtt{suc}(s)\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}=\emptyset\\ \bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}&\text{if }\mathtt{suc}(s)\neq\emptyset\text{ and }\mathrm{C}\neq\emptyset\end{cases}$
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model, and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\mathtt{suc}(s)$.
@@
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. Assume $\mathtt{suc}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\mathtt{suc}(s)\}$, which is a cover of $\mathtt{suc}(s)$. Assume $\mathtt{suc}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Let $t\in\mathtt{suc}(s)$. We want to show $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{nei}_{a}(s)$ is a cover of $\mathtt{suc}(s)$ for every $a\in\mathtt{AG}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in\mathrm{X}_{a_{1}}$, …, there is $\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in\mathrm{X}_{a_{n}}$. Then $t\in\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$. Note $\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
@@
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$, $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
@@
Alurz, Henzinger, and Kupferman [2] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Goranko and Jamroga [4] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.

### v3

@@ 4.2 Single-coalition-first action models
An action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is a single-coalition-first action model if the following conditions are met:
  • for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$: $\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$ is a general cover of $\mathsf{out}_{\emptyset}(s,\emptyset)$.
  • for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$, $s\in\mathtt{ST}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{\mathrm{C}}}|_{a})\mid a\in\mathrm{C}\}$.
It can be observed that for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathsf{out}_{\mathrm{C}}$ can be derived from $\mathsf{out}_{\emptyset}$ and all $\mathsf{out}_{a}$ where $a\in\mathrm{C}$. This is why we use “single-coalition-first”.
@@
For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ is equal to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
The following result offers two alternative definitions of single-coalition-first action models.
@@
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
@@
      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{\mathrm{C}}}|_{a})\mid a\in\mathrm{C}\}$.
@@
      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$: $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}\neq\emptyset$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. It suffices to show $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\emptyset}(s,\emptyset)$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $a\in\mathrm{D}$. By (1b), $t\in\mathsf{out}_{a}(s,\sigma_{\mathrm{D}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
First, we show (3a). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. It suffices to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
@@
By (1a) and (1b), it is easy to verify $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Assume $\overline{\mathrm{C}}$$\emptyset$. Clearly, what we want holds. Assume $\overline{\mathrm{C}}$$\neq\emptyset$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there is $\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})$, …, there is $\sigma_{a_{n}}\in\mathtt{JA}_{a_{n}}$ such that $w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
@@
Let $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}$. Then, there is $\sigma_{a}\in\mathtt{JA}_{a}$ such that $t\in\mathsf{out}_{a}(s,\sigma_{a})$. By (3a), there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{a}\subseteq\sigma_{\mathtt{AG}}$ and $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $t\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\emptyset\subseteq\sigma_{{\mathtt{AG}}}\}$, which is equal to $\mathsf{out}_{\emptyset}(s,\emptyset)$ by (3a).
@@
The first part of the result follows from the previous theorem. For the second part, the grand-coalition-first action model given in Example 1 is a counterexample.
@@
Goranko and Jamroga [goranko_comparing_2004] discussed convex concurrent game models, which are closely related to rectangular game forms in game theory [abdou_rectangularity_1998]. It can be verified that convex concurrent game models are single-coalition-first action $\mathtt{SID}$-models.
@@ 4.3 Single-coalition-first neighborhood models
Let $\mathtt{ST}$ be a nonempty set of states and $\Delta_{1},\Delta_{2}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}$. Define $\Delta_{1}\odot\Delta_{2}=\{\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\mid\mathrm{Y}_{1}\in\Delta_{1},\mathrm{Y}_{2}\in\Delta_{2},\text{and }\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset\}$. Let $I$ be a nonempty set of indices. Define $\bigodot\{\Delta_{i}\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\}$ as expected [^3] Strictly speaking, we should write $\bigodot\Big\{(i,\Delta_{i})\mid i\in I\text{ and }\Delta_{i}\subseteq\mathcal{P}(\mathtt{ST})-\{\emptyset\}\Big\}$. Here, we abuse the notation a bit.. Note $\bigodot\{\Delta\}=\Delta$.
@@
A neighborhood model $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},$ $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is a single-coalition-first neighborhood model if the following conditions are met:
  • for all $s\in\mathtt{ST}$, $\mathtt{nei}_{\emptyset}(s)$ is either empty or a singleton with a nonempty element.
  • for all $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{a}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
  • for every nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$: $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$.
The following fact will be used later.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}_{\mathrm{C}}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$.
@@
Let $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$. Assume $\bigcup\mathtt{nei}_{\emptyset}(s)=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\emptyset$, which is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$. Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}=\emptyset$. Then $\mathtt{nei}_{\mathrm{C}}(s)=\{\bigcup\mathtt{nei}_{\emptyset}(s)\}$, which is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$. Assume $\bigcup\mathtt{nei}_{\emptyset}(s)\neq\emptyset$ and $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$. Let $t\in\bigcup\mathtt{nei}_{\emptyset}(s)$. We want to show $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$. Note $\mathtt{nei}_{a}(s)$ is a cover of $\bigcup\mathtt{nei}_{\emptyset}(s)$ for every $a\in\mathtt{AG}$. Then, there is $\mathrm{X}_{a_{1}}\in\mathtt{nei}_{a_{1}}(s)$ such that $t\in\mathrm{X}_{a_{1}}$, …, there is $\mathrm{X}_{a_{n}}\in\mathtt{nei}_{a_{n}}(s)$ such that $t\in\mathrm{X}_{a_{n}}$. Then $t\in\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}$. Note $\mathrm{X}_{a_{1}}\cap\dots\cap\mathrm{X}_{a_{n}}\in\mathtt{nei}_{\mathrm{C}}(s)$. Then $t\in\bigcup\mathtt{nei}_{\mathrm{C}}(s)$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
@@
  • $\mathtt{S}\textbf{-}\mathtt{NM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\mathrm{Y}_{1}\in\mathtt{nei}_{\mathrm{C}}(s)$ and $\mathrm{Y}_{2}\in\mathtt{nei}_{\mathrm{D}}(s)$: $\mathrm{Y}_{1}\cap\mathrm{Y}_{2}\neq\emptyset$.
@@
Alur, Henzinger, and Kupferman [alur_alternating-time_1998] proposed a semantics for $\mathsf{ATL}$, where models are based on alternating transition systems. Goranko and Jamroga [goranko_comparing_2004] discussed tight alternating transition systems, which we can verify are single-coalition-first neighborhood $\mathtt{SID}$-models.